OpenStudy (anonymous):

Help me solve this exponential equation plz!! 256^(5x)=64^(x+4)

OpenStudy (anonymous):

4^4=256 and 4^3=64 so 256^(5x)=4^(4*5x) and 64^(x+4)=4^(3*(x+4)) u can substitue and solve from there

OpenStudy (anonymous):

2^8^(5x) = 2^6^(x+4) 40x=6x+24 34x=24 x=12/17

OpenStudy (anonymous):

I guess 4 would have been easier 20x=3x+12 17x=12 x=12/17 :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

it would still come out to 12/17 either method correct?

OpenStudy (anonymous):

yup as mrjoselson's steps showed. 4 would just be faster

OpenStudy (anonymous):

Thank you both =)